OpenStudy (anonymous):

(2sqrt3 +3)(3sqrt3-5)

OpenStudy (kainui):

FOIL means you multiply First, Outer, Inner, and Last together and then add them all up.

OpenStudy (anonymous):

I did that. I think I see what I did wrong though.. hold on

OpenStudy (anonymous):

3-sqrt3 ?

hero (hero):

(2√(3)+3)(3√(3)-5) = (2√(3)(3√(3)-5) + 3(3√(3)-5) = 6(3) - 10√3 +9√3 - 15 = 18 - 15 -√(3) = 3 - √(3)

OpenStudy (anonymous):

Yay, Thanks guys. I have another question I'm unsure about.

OpenStudy (anonymous):

27^-(2/3)

hero (hero):

\[27^{-2/3} = \sqrt[3]{27^{-2}}= \sqrt[3]{\frac{1}{27^2}}=\sqrt[3]{\frac{1}{9^{(3\dot\ 2)}}}=\sqrt[3]{\frac{1}{9^{(2 \dot\ 3)}}}=\sqrt[3]{\frac{1^3}{81^3}}=\sqrt[3]{\left(\frac{1}{81}\right)^3}=\frac{1}{81}\]

hero (hero):

Just because you close a question, that doesn't mean it is over. You can still communicate here if needed.

OpenStudy (agent0smith):

@Hero you made a mistake, it should be 1/9 not 1/81. \[27^{2} = 3^{3 \times 2}\]not \[9^{2 \times 3}\]